OpenStudy (anonymous):

00:53:28 Mrs. Grijalva hired a landscaping company to landscape her back yard. The company has a standing policy to plant 3 bushes for every tree planted. The company charges $75 for each tree planted and $20 for each bush. Mrs. Grijalva was billed for $810 after the planting was complete. How many bushes did the company plant in Mrs. Grijalva’s back yard? A. 2 B. 6 C. 18 D. 54

OpenStudy (anonymous):

We get 3x as many bushes as trees. B=3T Bushes cost $20 each, and trees costs $75 each Cost=$20∗B+$75∗T Total cost was $810 810=20B+75T Now solve by substitution, writing 3T wherever we see B in the cost equation: 810=20(3T)+75T 810=135T T=6 Plug into formula giving # of bushes for number of trees B=3T=3(6)=18

OpenStudy (anonymous):

Hope I helped